# The expected value of sample information calculations for external validation of risk prediction models

### v2

@@ Introduction
Clinical prediction models that quantify the risk of medically significant events are a key aspect of individualized medicine. Before being trusted for clinical use in a target population, a model needs to undergo validation in a representative sample from that population. Model performance during such external validation is assessed in various ways, which can broadly be classified as prediction error (e.g., Brier score), discrimination (e.g., c-statistic), calibration (e.g., calibration intercept and slope), and net benefit (NB)1. Among such metrics, NB is a decision-theoretic one, as it enables direct assessment of the clinical utility of a model (i.e., if a model’s expected NB is higher than alternatives, it is expected to confer clinical utility). Due to such strong decision-theoretic underpinnings, and despite its relatively new arrival in the field of predictive analytics, NB has gained significant momentum and has become a standard component of modern external validation studies2.
@@
When interpreting the results of a validation study, the finite size of the validation sample means that the assessment of model performance is accompanied by uncertainty. Uncertainty in conventional metrics of model performance is typically communicated using classical inferential methods (e.g., 95% confidence interval [CI] around c-statistic or calibration slope, p-values for comparing the c-statistic of competing models). However, the relevance of such frequentist approach to uncertainty around NB as a decision-theoretic metric is doubtful3–5.
@@
Decision theory provides an alternative view to the consequences of uncertainty by relating it to the outcome of decisions. Leaving aside, for the sake of simplicity, the issue of risk aversion, the decision-maker should adopt the strategy with the highest expected utility, which, if implementation costs are comparable, is the one with the highest net benefit. While uncertainty around NB should not affect the ‘adoption decision’, it is associated with utility loss as it hinders our ability to identify the decision with the highest true NB. The extent of uncertainty should thus inform the ‘research decision’: whether the expected utility loss is large enough to necessitate collecting further evidence. This approach towards uncertainty quantification is referred to as Value of Information (VoI) analysis6–8.
@@
VoI methods are widely accepted in health policy-making7,9. However, their relevance to clinical decision-making and risk prediction has only recently been highlighted10. In particular, the expected value of perfect information (EVPI), the expected gain in NB by completely eliminating uncertainty, has been applied to both the development and validation phases of risk prediction models10,11. If based on EVPI, the investigator concludes that uncertainty is large and procuring more validation data is warranted, the next natural question is ‘how much data should be collected?’. The expected value of sample information (EVSI) is the expected gain in NB by conducting a study of a given sample size (and design)12. In the context of model validation, EVSI can thus inform the sample size determination of future validation studies.
@@
The aim of the present work is to define EVSI for the external validation phase of a risk prediction model and propose algorithms for its computation. The rest of this manuscript is structured as follows. After outlining the context, we define EVSI for a future study aimed at evaluating the clinical utility of a model in a target population. We propose algorithms for EVSI computation based on bootstrapping when preliminary data from the target populations are available, a beta-binomial modeling approach for binary outcomes when current information can be specified as independent Beta distributions, and a more general algorithm when the current information is expressed as an arbitrary probability distribution. A case study puts the developments in context where we provide practical suggestions on reporting and interpreting EVSI. We conclude by proposing areas of further inquiry.
@@ Methods
We focus on validating a previously-developed risk prediction model in a well-defined new target population. The developments are presented for models for binary responses but are generally applicable to any context where NB can be assessed (e.g., survival outcomes13, models for treatment benefit14). The model is advertised as a function that maps patient characteristics to an estimate of event risk. We acknowledge the relevance of other sources of uncertainty (e.g., missing data, the plausibility that the sample is truly representative of the target population) but focus the developments on sampling uncertainty, and discuss some other uncertainty sources whenever relevant.
VoI analysis is a Bayesian paradigm and requires explicit specification of current information on the performance of the model in the target population. By default, we assume this information is generated from a preliminary validation exercise based on an iid sample $\mathbf{d}$ of $n$ individuals. However, such information can be generated in other ways. For example, one can extrapolate (optimism-corrected) performance of a model from its development sample, or its performance in previous validation studies15, and accommodate any added uncertainty about the compatibility of the populations (e.g. by quantitatively discounting external information16). Formal expert elicitation approaches can also be employed to construct distributions around model performance metrics17. In such instances, $\mathbf{d}$ can be an abstract entity representing, for example, the expert-elicited information.
Given our current information, we are considering the utility of procuring data of sample of size $n^{*}$ for a future validation study in the same target population. We are interested in expected utility gain from such a study. We express utility in terms of net benefit (NB), as is quantified and presented in decision curve analysis18. In a nutshell, NB is based on the specification of a risk threshold, denoted by $z$, placed on predicted risks such that individuals with predicted risk lower than this threshold are considered low-risk, and those with predicted risk higher than this threshold are classified as high-risk (and therefore are selected for treatment). For individuals with predicted risk precisely equal to $z$, the decision-maker will be ambivalent about the treatment decision. Because among these individuals, treating all means a proportion $z$ is properly treated (true positive) while the proportion $1-z$ is unnecessarily treated (false positive), such ambivalence implies that each false positive has $-z/(1-z)$ times clinical utility of a true positive. Because the true and false positive rates of the model can be estimated at this threshold, this implied exchange rate can be used to evaluate the NB of a model in net true positive or false positive units. Any given model should be compared with at least two ‘default’ strategies of treating no one and treating all. Treating no one has NB=0 by definition, and NB calculations for treating all follow the same logic as above, but by considering all individuals as positive. In practice, NB is measured over a range of plausible risk thresholds.
Let $\theta$ be the set of parameters that fully specify the target population for NB calculations. A common setup for binary outcomes is to define $\theta$ as the triplet of prevalence ($\theta_{p}$), sensitivity ($\theta_{se}$), and specificity ($\theta_{sp}$)20. We index treating no one, model-based treatment, and treating all by, respectively, subscripts 0,1,2, resulting in the following equation for NBs (for brevity, we drop the notation that indicates quantities are functions of $z$)
@@
with $\theta=\{\theta_{p},\theta_{se},\theta_{sp}\}$. For time-to-event outcomes, these quantities are time-dependent and should be defined at a time-horizon of interest13. If there are alternative strategies they can also be considered (the above indexing of $NB()$ can accommodate other strategies), but this is omitted from this work for the ease of exposition.
Our current information about $\theta$ based on the available evidence is expressed as $P(\theta|\mathbf{d})$. According to Bayes’ theorem, $P(\theta|\mathbf{d})\propto P(\theta)P(\mathbf{d}|\theta)$, indicating that this information is influenced by any prior information we might have had, as well as by what we have learned from $\mathbf{d}$.
@@
With $P(\theta|\mathbf{d})$ quantifying their knowledge about $\theta$, a risk-neutral decision-maker should pick the strategy with the highest expected NB given the current information:
$\mathbb{E}NB_{\text{current information}}=\max_{i}{\mathbb{E}_{\mathbb{\theta}|\mathbf{d}}NB(i,\mathbb{\theta})},$ (2)
where the expectation is with respect to $P(\theta|\mathbf{d})$. Generally, evaluating this expectation might require a Monte Carlo approach (e.g., via Markov Chain Monte Carlo sampling from $P(\theta|\mathbf{d})$). However, in many instances, this expectation can be more readily obtained. For example, a Bayesian interpretation of the bootstrap allows one to interpret an estimate of $\theta$ obtained from a bootstrapped copy of $\mathbf{d}$ as a random draw from the posterior distribution of $\theta$ given $\mathbf{d}$21. However, as the number of bootstrap iterations increases, due to the law of large numbers the average values from such bootstrap samples for each component of $\theta$ as well as $NB$s converge to their point estimate in the sample. This justifies interpreting the point estimates of $NB$ from the sample, as is conducted in conventional decision curve analysis21, in a Bayesian framework as the posterior mean under a non-informative prior.
@@
Details of the reasoning and algorithms behind calculating EVPI is presented previously11. In brief, if we could know the true values of $\theta$, we would have picked the strategy with the highest true NB. We indeed do not have access to the ground truth, but have partial information in terms of $P(\theta|\mathbf{d})$. As such, we can model the consequence of knowing the truth and take its expectation with respect to this distribution:
$\mathbb{E}NB_{\text{perfect information}}=\mathbb{E}_{\mathbf{\theta}|\mathbf{d}}\max_{i}NB(i,\mathbf{\theta}).$ (3)
@@
$EVPI=\mathbb{E}NB_{\text{perfect information}}-\mathbb{E}NB_{\text{current information}}.$ (4)
The EVPI is a scalar quantity in NB units and quantifies the expected loss in NB due to uncertainty, or alternatively, the expected gain in NB by completely resolving uncertainty (i.e., conducting a validation study of infinite size). We have proposed a generic algorithm based on bootstrapping to estimate this expectation, as well as an asymptotic approach based on central limit theorem and two-dimensional unit normal loss integral11.
@@
The reasoning behind EVSI calculation is similar to that of EVPI, with the modification that instead of knowing the truth, we will have more (but not complete) information about $\theta$, with additional information coming from a future iid sample $\mathcal{D^{*}}$ of size $n^{*}$. With this added information we will have the opportunity to revise our current decision. We will update our knowledge from $P(\theta|\mathbf{d})$ to $P(\theta|\mathbf{d},\mathcal{D^{*}})$, and will select the strategy that has the highest expected $NB$.
The NB of this approach once future data $\mathcal{D^{*}}$ are obtained is $\max_{i}\mathbb{E}_{\mathbf{\theta|{\mathcal{D^{*}},d}}}NB(i,\mathbf{\theta})$. We do not know what data we will observe in the future, but we can treat $\mathcal{D^{*}}$ as a random variable, and take the expectation of this line of reasoning with regard to its predictive distribution:
$\mathbb{E}NB_{\text{sample information}}=\mathbb{E}_{\mathcal{D^{*}}|\mathbf{d}}\max_{i}\mathbb{E}_{\mathbf{\theta}|\mathcal{D^{*}},\mathbf{d}}NB(i,\mathbb{\theta}).$ (5)
@@
$EVSI=\mathbb{E}NB_{\text{sample information}}-\mathbb{E}NB_{\text{current information}}.$ (6)
The EVSI is a non-negative scalar quantity in the same NB units as EVPI. It quantifies the expected gain in NB by procuring a further validation sample of a given size. The higher the EVSI, the higher the expected gain from the planned validation study. EVPI, quantifying the expected NB gain by completely resolving uncertainty, puts an upper limit on EVSI.
@@
The challenge for EVSI calculation is the two nested expectations. The inner expectation represents updated estimates of NBs after obtaining future data, and the outer expectation is with respect to the predictive distribution of future data. Below we detail three computaion algorithms.
@@
If the current evidence $\mathbf{d}$ is in terms of a random sample of size $n$ from the target population of interest, sampling from $P(\theta|\mathbf{d})$ can be performed via the Bayesian bootstrap21. The resulting algorithm for EVSI computation is similar to the previously proposed method for EVSI calculations for data-driven economic evaluations22 and can be seen as an extension of the bootstrap method for EVPI computation for external validation studies11.
The Bayesian bootstrap of $\mathbf{d}$ assigns an improper non-informative prior distribution $\text{Dirichlet}(0,0,...,0)$ across the $n$ data points. A random draw from the posterior distribution of the generating population can be materialized by drawing a random weight vector from the $\text{Dirichlet}(1,1,...,1)$ distribution. Estimating $\theta$ in this weighted sample is akin to drawing a value for $\theta$ from its posterior distribution21. Further, once the generating population is at hand, the data of the future study ($\mathcal{D^{*}}$) can be drawn by sampling from this population. This results in a two-level resampling approach for EVSI calculation explained in Table 1.
[…]
${}^{b}$ $\mathbb{E}NB_{\text{current information}}$ can also be directly evaluated in the original sample without bootstrapping. However, we recommend averaging over bootstrapped values as this approach is consistent with the way other terms are evaluated and this prevents getting occasional negative VoI metrics

### v3

@@ Introduction
Risk prediction models that quantify the probability of clinical events are a key aspect of individualized medicine. Once developed, these models are often used in diverse populations. Before being trusted for clinical use in a new population, a prediction model needs to undergo validation in a representative sample from that population1. For example, the original Framingham risk scores for cardiovascular disease risk are developed using data from the Framingham County in the US, but are externally validated for use in many different populations2.
@@
During a validation study, the performance of the prediction model is assessed in terms of average prediction error, discrimination (e.g., c-statistic), calibration (e.g., calibration intercept and slope), and net benefit (NB)3. Among such metrics, NB is a decision-theoretic one, as it enables direct assessment of the clinical utility of a model (i.e., if a model’s expected NB is higher than that of alternatives, it is expected to confer clinical utility). Due to such decision-theoretic underpinnings, and despite its relatively new arrival in the field of predictive analytics, NB has gained significant momentum and has become a standard component of modern external validation studies4.
@@
When interpreting the results of a validation study, the finite size of the validation sample means that the assessment of model performance is accompanied by uncertainty. Uncertainty in conventional metrics of model performance is communicated using classical inferential methods (e.g., 95% confidence interval [CI] around c-statistic or calibration slope). Similarly, sample size considerations when planning a validation study are based on inferential statistics5,6. However, the relevance of this approach to uncertainty around NB as a decision-theoretic metric is doubtful, as arbitrary significance levels or pre-specified confidence bands do not necessarily translate to better decisions7–9.
@@
Decision theory provides an alternative view to the consequences of uncertainty by relating it to the outcome of decisions. A decision-maker should adopt the strategy with the highest expected utility, which, if implementation costs are comparable, is the one with the highest expected NB. While uncertainty around NB should not affect the ‘adoption decision’, it is associated with utility loss as it hinders our ability to identify the decision with the highest true NB. The extent of uncertainty should thus inform the ‘research decision’: whether the expected utility loss is large enough to necessitate collecting further evidence. This approach towards uncertainty quantification is referred to as Value of Information (VoI) analysis10–12.
@@
VoI methods are widely accepted in health policy-making11,13. However, their relevance to clinical decision-making and risk prediction has only recently been highlighted14. In particular, the expected value of perfect information (EVPI), the expected gain in NB by completely eliminating uncertainty, has been applied to both the development and validation phases of risk prediction models14,15. However, to the best of our knowledge, the expected value of sample information (EVSI), the expected gain in NB by conducting a study of a given sample size, has not been defined and applied to clinical prediction models.
@@
The aim of the present work is to define EVSI for the external validation studies of risk prediction models and propose algorithms for its computation. The rest of this manuscript is structured as follows. After outlining the context, we define EVSI for a future validation study aimed at evaluating the clinical utility of a model in a new population. We propose different algorithms for EVSI computations, elaborate on their use case, and compare their performance in simulation studies. A case study puts the developments in context. We conclude by proposing areas of further inquiry.
@@ Methods
We focus on validating a previously-developed risk prediction model in a new target population. The developments are presented for prediction models for binary responses but are generally applicable to any context where NB can be assessed (e.g., survival outcomes16, models for treatment benefit17). A risk prediction model is advertised as a function that maps patient characteristics to an estimate of event risk. We have some ‘current’ information about the performance of this model in the target population (for example, from expert opinion, or from a pilot study), but are uncertain if using the model in this population is net beneficial. We are planning to obtain a random sample $\mathcal{D}$ containing $\mathit{N}$ observations from this population to reduce such uncertainty. This sample is expected to improve our knowledge about model performance, which in turn will increase our chance of making the correct decision whether to use the model in this population. As such, procuring such a sample might be associated with NB gain, and we are interested in evaluating the expected gain in NB as a function of $\mathit{N}$. We acknowledge the relevance of other sources of uncertainty (e.g., missing data, the plausibility that the sample is truly representative of the target population) but focus on sampling uncertainty – that is, our uncertainty in true NBs because of the finite size of the validation sample.
Net benefit calculations for risk prediction models
Details of NB calculations for risk prediction models are explained in its original paper and in many tutorials4,18,19. In brief, to turn a continuous predicted risk to a binary classification to inform a treatment decision, a decision-maker needs to specify a risk threshold $z$ on predicted risks such that individuals classified as high-risk are selected for the treatment (e.g., a 7.5%-10% threshold on Framingham risk score is often used to initiate treatment with statins to reduce cardiovascular disease risk20). Compared to treating no one (with a default utility of 0), the utility of this classification can be expressed as $p_{tp}-\omega p_{fp}$, where $p_{tp}$ and $p_{fp}$ are the probabilities of true positive and false positives, respectively, at this threshold, and $\omega\geq 0$ represents the utility tradeoff (exchange rate) between a true positive and a false positive classification. Vickers and Elkin showed that the value of $\omega$ can be deduced from the risk threshold itself18: if a decision-maker is willing to treat those with risk $>z$ and not treat those with risk $<z$, it means among individuals with risk precisely equal to $z$, they will be ambivalent about the treatment decision. This ambivalence indicates that in a population whose average outcome risk is $z$, the expected utility of treating everyone and treating no one is the same to the decision-maker. Because in this population, treating everyone corresponds to $p_{tp}=z$ and $p_{fp}=1-z$, we have $\omega=z/(1-z)$. For example, a 10% threshold on cardiovascular risk for initiating statin therapy means the benefit of treating one individual who will experience a cardiovascular event is considered equal to the harm of treating nine individuals who will not experience such an event without treatment. This exchange rate enables calculating NB in either net true positive or net false positive units for a given threshold. In practice, the NB is calculated across a plausible range of thresholds.
It is often more intuitive to express $p_{tp}$ and $p_{fp}$ in terms of outcome prevalence and the model’s true sensitivity and specificity in the target population21. Let $\theta$ be the triplet of outcome prevalence ($\theta_{p}$, a quantity that is independent of model performance), sensitivity ($\theta_{se}$), and specificity ($\theta_{sp}$). We have $p_{tp}=\theta_{p}\theta_{se}$ and $p_{fp}=(1-\theta_{p})(1-\theta_{sp})$ (for brevity, we drop the notation that indicates sensitivity and specificity are functions of $z$).
Any given model should be compared with at least two ‘default’ strategies of treating no one and treating all. Treating no one has NB=0 by definition, and NB calculations for treating all follow the same logic as above, but by considering all individuals as positive. Thus we have three competing strategies: do not treat anyone, use the model to treat those with predicted risk $\geq z$, and treat all. We index them by 0,1,2, respectively. This results in the following equation for NBs:
@@
For time-to-event outcomes, $\theta$s and thus NBs are time-dependent and should be defined at a time-horizon of interest16. If there are alternative strategies, they can also be considered (the above indexing of NB() can accommodate other strategies), but this is omitted here for the ease of exposition.
Value of Information analysis
VoI analysis is a Bayesian approach and requires explicit specification of current information on the performance of the model in the target population. By default, we assume this information is generated from a preliminary validation exercise based on a random sample $\mathbf{d}$ of size $n$. However, such information can be generated in other ways. For example, one can extrapolate the performance of a model from its development sample, or its performance in previous validation studies22, and accommodate any added uncertainty about the compatibility of the populations by quantitatively discounting external information23. Formal expert elicitation approaches can also be employed to construct distributions around model performance metrics24. In such instances, $\mathbf{d}$ can be an abstract entity representing, for example, the expert-elicited information.
@@
Our current information about $\theta$ based on the available evidence is expressed as $P(\theta|\mathbf{d})$. According to Bayes’ theorem, $P(\theta|\mathbf{d})\propto P(\theta)P(\mathbf{d}|\theta)$, indicating that this information is influenced by any prior information we might have had, as well as by $P(\mathbf{d}|\theta)$, i.e., the likelihood function, the support the observed data provide for a given value of $\theta$. With such current information, the decision-maker should pick the strategy with the highest expected NB given the current information:
${{E}}NB_{\text{current}}=\max_{{}_{i\in\{0,1,2\}}}{\mathbb{E}_{\mathbb{\theta}|\mathbf{d}}NB(i,\mathbb{\theta})},$ (2)
where the expectation is with respect to $P(\theta|\mathbf{d})$.
@@
Details of the reasoning and algorithms behind calculating EVPI are presented previously15. In brief, if we could know the true values of $\theta$, we would pick the strategy with the highest true NB. We indeed do not have access to the ground truth, but have partial information in terms of $P(\theta|\mathbf{d})$. As such, we can model the consequence of knowing the truth (having perfect information) and take its expectation with respect to this distribution:
${{E}}NB_{\text{truth}}=\mathbb{E}_{\mathbf{\theta}|\mathbf{d}}\left[\max_{{}_{i\in\{0,1,2\}}}NB(i,\mathbf{\theta})\right].$ (3)
@@
$EVPI={{E}}NB_{\text{truth}}-{{E}}NB_{\text{current}}.$ (4)
The EVPI is a scalar quantity and quantifies the expected loss in NB due to uncertainty, or alternatively, the expected gain in NB by completely resolving uncertainty (i.e., conducting a validation study of infinite size). We have proposed a generic algorithm based on bootstrapping to estimate this expectation, as well as an asymptotic approach based on central limit theorem and two-dimensional unit normal loss integral15.
@@
The reasoning behind EVSI calculation is similar to that of EVPI, with the modification that instead of knowing the truth, we will have more (but not perfect) information about $\theta$, with additional information coming from obtaining a future sample $\mathcal{D}$. With this added information we will have the opportunity to revise our current decision. We will update our knowledge from $P(\theta|\mathbf{d})$ to $P(\theta|\mathbf{d},\mathcal{D})$, and will select the strategy that has the highest expected NB.
The NB of this approach once future data are obtained is $\max_{i}\mathbb{E}_{\mathbf{\theta|{\mathcal{D},d}}}NB(i,\mathbf{\theta})$. We do not know what data we will observe in the future, but we can treat $\mathcal{D}$ as a random variable, and take the expectation of this line of reasoning with regard to its predictive distribution (the distribution for future observations given current data and the model):
${{E}}NB_{\text{sample}}=\mathbb{E}_{\mathcal{D}|\mathbf{d}}\left[\max_{{}_{i\in\{0,1,2\}}}\mathbb{E}_{\mathbf{\theta}|\mathbf{d},\mathcal{D}}NB(i,\mathbb{\theta})\right].$ (5)
@@
$EVSI={{E}}NB_{\text{sample}}-{{E}}NB_{\text{current}}.$ (6)
The EVSI is a non-negative scalar quantity in the same NB units as EVPI. It quantifies the expected gain in NB by procuring a validation sample of a given size. The higher the EVSI, the higher the expected gain from the planned validation study. EVPI, quantifying the expected NB gain by completely resolving uncertainty, puts an upper limit on EVSI.
@@
The challenge for EVSI calculation is the two nested expectations in ${{E}}NB_{\text{sample}}$. The inner expectation represents updated estimates of NBs after obtaining future data, and the outer expectation is with respect to the predictive distribution of future data. Below we detail three computation algorithms.
@@
If the current evidence $\mathbf{d}$ is in terms of a random sample of size $n$ from the target population (e.g., from a pilot study), EVSI computation can be performed via a two-level resampling algorithm. This approach has been previously applied to EVSI computation for data-driven economic evaluations25 and can be seen as a direct extension of the bootstrap method for EVPI computation for external validation studies15. The logic behind this approach is provided in detail previously25. In a nutshell, let $\mathcal{D^{*}}$ be the target population from which $\mathbf{d}$ and $\mathcal{D}$ are sampled. We take $\mathcal{D^{*}}$ as a random entity and apply the Bayes’s rule: $P(\mathcal{D^{*}}|\mathbf{d})\propto P(\mathcal{D^{*}})P(\mathbf{d}|\mathcal{D^{*}})$. Rubin showed that by using the improper prior $P(\mathcal{D^{*}})\sim\text{Dirichlet}(0,0,…,0)$ placed on all possible observations in $\mathcal{D^{*}}$, the posterior distribution $P(\mathcal{D^{*}}|\mathbf{d})$ will be a discrete distribution that puts random weights Dirichlet(1,1,…,1) on each observation in $\mathbf{d}$. Since any parameter of interest $\theta$ is a function of $\mathcal{D^{*}}$, posterior draws of weights from the $\text{Dirichlet}(1,1,…,1)$ distribution produce draws from the posterior distribution of $P(\theta|\mathbf{d})$26. Further, once the generating population is at hand, the data of the future study ($\mathcal{D}$) can be drawn by sampling from this population. The resulting two-level resampling algorithm is explained in Table 1.
- Prerequisite: Obtain $\mathbf{d}$, a pilot sample of size $n$ consisting of predicted risks and observed responses from the target population.
1. For $j$ = 1 to $M$ (number of Monte Carlo simulations)
(a) Draw $\mathcal{D^{*}}$, a Bayesian bootstrap of the data $\mathbf{d}$, as a random draw from the posterior distribution of the population that has generated the sample. This is done by assigning random weights $W^{*}\sim\text{Dirichlet}(1,1,...,1)$ to each observation in $\mathbf{d}$.
(b) Estimate $\theta^{*}$, the true values of prevalence, sensitivity, and specificity in this iteration, from $\mathcal{D^{*}}$.
(c) Calculate $NB^{*}_{0}$, $NB^{*}_{1}$, and $NB^{*}_{2}$, by plugging $\theta^{*}$ into Equation 1. These are true NBs in this iteration. Record the maximum of true NBs: $NB^{*}_{\text{truth}}=\max\{NB^{*}_{0},NB^{*}_{1},NB^{*}_{2}\}$.
(d) Draw $\mathcal{D}$, a random realization of the data of the future study, through sampling $\mathit{N}$ observations with replacement from $\mathcal{D^{*}}$.
(e) Create $\mathcal{D^{+}}=\mathbf{d}+\mathcal{D}$, the pooled data of existing and future samples.
(f) Calculate $\theta^{+}$ from $\mathcal{D^{+}}$, the updated estimates of prevalence, sensitivity, and specificity after obtaining the future sample.
(g) Calculate $NB^{+}_{0}$, $NB^{+}_{1}$, and $NB^{+}_{2}$, updated estimates of expected NBs after obtaining the future sample, by plugging $\theta^{+}$ into Equation 1. Determine the strategy that has the highest expected future NB: $l={\mbox{argmax}_{i\in\{0,1,2\}}(NB^{+}_{i})}$. Record the true NB of this strategy: $NB^{*}_{\text{sample}}=NB^{*}_{l}$.
- Next $j$
2. Take the average of $NB^{*}$s from step 1c and pick the maximum value. This is ${{E}}NB_{\text{current}}$§.
3. Average $NB^{*}_{\text{truth}}$s from step 1c. This is ${{E}}NB_{\text{truth}}$. From this subtract ${{E}}NB_{\text{current}}$. This is EVPI.
4. Average $NB^{*}_{\text{sample}}$s from step 1g. This is ${{E}}NB_{\text{sample}}$. From this subtract ${{E}}NB_{\text{current}}$. This is EVSI.
Ordinary bootstrap can also be used - see text.
One way of generating such wights is by generating $n$ exponential(1) random variables and scaling them to add up to 1.
§ Because the NB terms are linear on the elements of $\theta$, ${{E}}NB_{\text{current}}$ can also be directly evaluated in the original sample without bootstrapping, by plugging in the point estimates of $\theta$ into Equation 1. However, we recommend averaging over bootstrapped values as the positive correlation between this term and the other term in EVPI and EVSI equations improves precision and prevents getting occasional negative VoI estimates.